OpenStudy (anonymous):

Can somone help me Let f(x)=2x+2,g(x)=x^2 ,and h(x)=3/x-2. State the domain od the operation. h(g(x))

OpenStudy (anonymous):

\[h(g(x))=h(x^2)=\frac{3}{x^2-2}\] you may not divide by zero, so set \[x^2-2=0\] solve for x to get \[x=\pm \sqrt{2}\] so domain is all real numbers except \[x=\pm \sqrt{2}\]